# How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching?

### v6

@@ 5.5. Evaluation & Ablation Study
Figure 11 compares the matching performance with/without using the pipeline approaches: ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of the above (CR). We can see that ad hoc logic-based repair (AR) and post hoc LLM-based repair (PR) can significantly improve precision, with only a slight decrease in recall, but the overall F1 score also increases. Although there is no restriction on the use of both ad hoc and post hoc repairs, the combination does not improve precision, recall, or F1 score. For syntactic OM, its main purpose is to provide “seed mappings” for lexical and semantic matching. These mappings are meant to be precise and accurate. Both of our pipeline repair approaches have shown high precision and are the best tool suites.
@@
Figure 11. Evaluation of pipeline repair approaches: Ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of above (CR). The symbol $\varnothing$ represents the baseline without text preprocessing.
Figure 12 compares post hoc LLM-based repair (PR) with two alternative repair approaches. The results show that PR achieves almost the same precision with text preprocessing only repair (ToR) and LLM only repair (LoR), but its recall and F1 score are better than the other two alternatives. For syntactic OM, this means that our approach could detect more ”seed mappings”.
@@ 6. Conclusion
OM aims to bridge the gap by capturing similar concepts that occur in different ontologies. OM is a complex process that combines syntactic matching, lexical matching, and semantic matching. Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching. This study concentrates on the common practice of using the text preprocessing pipeline in syntactic OM. While generic text preprocessing pipeline methods in syntactic OM are usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
@@
    This study demonstrates that the nature of text preprocessing used in NLP applications can be context-based. Not all pipeline methods are effective, and some of them may even hamper the performance of downstream tasks. Modifications are required before applying text preprocessing to specific NLP tasks.

### v7

@@ 5.5. Evaluation
Figure 12(c) compares the matching performance with/without using the pipeline approaches: ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of the above (CR). We can see that ad hoc logic-based repair (AR) and post hoc LLM-based repair (PR) can significantly improve precision, with only a slight decrease in recall, but the overall F1 score also increases. Although there is no restriction on the use of both ad hoc and post hoc repairs, the combination does not improve precision, recall, or F1 score. For syntactic OM, its main purpose is to provide “seed mappings” for lexical and semantic matching. These mappings are meant to be precise and accurate. Both of our pipeline repair approaches have shown high precision and are the best tool suites.
@@
Figure 12. Evaluation of pipeline repair approaches: Ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of above (CR). The symbol $\varnothing$ represents the baseline without text preprocessing repair.
[…]
Evaluation of pipeline repair approaches.
Figure 13 evaluates the scalability of ad hoc logic-based repair (AR). The x-axis indicates the number of entities in each alignment, and the y-axis shows the number of reserved words per 100 entities. For all the alignments we tested, the value of reserved words @100 is less than 10. With the number of entities increasing (i.e. the ontology size increases), the value of reserved words @100 does not increase but decreases. This feature of our approach is scalable in large-scale syntactic OM.
Figure 13. Comparison of reserved words per 100 entities (@100) with the number of entities in each alignment.
[…]
Comparison of reserved words per 100 entities (@100) with the number of entities in each alignment.
Figure 14(c) compares post hoc LLM-based repair (PR) with two alternative repair approaches. The results show that PR achieves almost the same precision with text preprocessing only repair (TOR) and LLM only repair (LOR), but its recall and F1 score are better than the other two alternatives. For syntactic OM, this means that our approach could detect more ”seed mappings”.
@@
[…]
Comparison of post hoc LLM-based repair (PR) with two alternative repair approaches.
@@ 6. Conclusion
While the classic text preprocessing pipeline in syntactic OM is usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
@@
    We showcase the limitations and opportunities for integrating the text preprocessing pipeline with LLMs in syntactic OM. We find that LLMs are good tools to find false mappings, but their outputs for finding true mappings are non-deterministic. The best practice is to integrate the classic text preprocessing pipelines with the LLMs. Using function calling to invoke the classic text preprocessing pipeline within LLMs, our approach could produce more stable and robust “seed mappings” than traditional text preprocessing methods and purely prompt-based approaches.
  4. (4)
    Our study also demonstrates that the nature of text preprocessing used in NLP applications can be context-based. Not all pipeline methods are effective, and some of them may even hamper the performance of downstream tasks. Modifications are required before applying text preprocessing to specific NLP tasks.
@@ Acknowledgements.
Appendix A Missing References in the LargeBio Track
As mentioned in the Limitations Section, the OAEI tracks are undergoing continuous development and may miss some reference alignments. Here we list the missing reference alignments found by the text preprocessing pipeline in the folder “csvs”. These mappings have outstanding possibilities of being correct, and we consider these mappings to be TPs in our study.